OpenStudy (eturpin1):

solve algebraically and prove if it is or is not an identity secx+cscx=1

OpenStudy (anonymous):

\[\begin{align*} \sec x+\csc x&=\frac{1}{\cos x}+\frac{1}{\sin x}\\\\ &=\frac{\sin x}{\sin x\cos x}+\frac{\cos x}{\sin x \cos x}\\\\ &=\frac{\sin x+\cos x}{\sin x\cos x} \end{align*}\] For the right hand side to simplify to \(1\), you need to show that \[\sin x+\cos x=\sin x\cos x\] Is this true?

OpenStudy (eturpin1):

No, it is not